# Driving Animatronic Robot Facial Expression From Speech

### v2

@@ V Skinning Motion Imitation Learning
Figure 5: Motion Space Validation. Actuated blendshape error for different facial regions (left figure): Color-coded skinning landmarks represent different facial regions for evaluation. MSE error distributions between simulated and physically actuated blendshapes are presented using violin plots, box and whisker plots, and scattered points, with each point representing one blendshape. Blendshapes are grouped by facial region, and only landmarks in the corresponding region are used for evaluation. Median errors are $2.41\mathrm{mm}$ (eye), $3.27\mathrm{mm}$ (brow), $1.76\mathrm{mm}$ (nose), $4.01\mathrm{mm}$ (cheek), $3.76\mathrm{mm}$ (mouth), and $8.63\mathrm{mm}$ (jaw). Qualitative comparison (right figure): Eight comparisons between simulated and actuated blendshapes are shown. Blendshapes (1)-(6) demonstrate high accuracy, while (7) mouth close and (8) jaw open highlight limitations in the current design, exhibiting maximum errors for their respective regions.
@@ V-A Model Architecture
The model architecture $f(\cdot)$ takes raw speech waveforms $\mathbf{s}$ and generates blendshape coefficients $\boldsymbol{\theta}$ representing skinning motions. It contains: (i) a frame-level speech encoder extracting embeddings via a transformer-based phoneme logit extractor and temporal fusion module, (ii) a speaking style encoder embedding conditioning vectors, and (iii) an LBS encoder projecting conditioned speech embeddings to blendshape coefficients in the range $[0,1]$.
Specifically, the phoneme extractor follows the state-of-the-art transformer-based self-supervised pre-trained speech model, Wav2vec2 [31], finetuned on the CommonVoice dataset with $53$ languages and $392$ phoneme classes to enhance cross-language generalization. It outputs a phoneme logits sequence at $49\mathrm{Hz}$ followed by resampling to $25\mathrm{Hz}$ to match the robot orchestration rate. The temporal fusion module progressively fuses $K$-step neighborhood logits over a sliding window into frame-level speech embeddings, comprising $M=\log_{2}K$ stacked Resnet blocks with a kernel size of $3$, stride of $2$, and $H$ filters. The $N$-way speaking style vector accounts for cross-subject variations and is embedded with a fully-connected (FC) layer having $H$ hidden units. The LBS encoder takes the late-fused conditioned speech embedding as input and outputs blendshape coefficients with two successive layers: an FC layer having $2H$ hidden units with a ReLU activation, and an FC layer with $H$ hidden units followed by a Sigmoid activation.
@@ V-C Inference
The robot LBS decoder transforms blendshape coefficients into skinning motions using designed robot blendshapes with frozen weights. At inference, the trained model first generates a sequence of blendshape coefficients from input speech. A low-pass Butterworth filter then smooths the sequence over time for temporal stability. The filtered sequence is input to the robot LBS decoder to compute the skinning motion as the reference signal in the proposed kinematics simulator.
@@ V-D Implementation Details
The VOCASET dataset [26] of audio-3D scan pairs of English utterances is used for training and testing. VOCASET was chosen for its high-quality 3D facial motion captures, diverse subjects, and phonetically balanced utterances, making it well-suited for learning speech-driven facial animations. The dataset contains $255$ unique sentences, some shared across $12$ subjects. VOCASET has $480$ facial motion sequences captured at $60\mathrm{Hz}$, each lasts $3$ to $4$ seconds. The 3D face mesh has $V=5023$ vertices. For our purposes, VOCASET motions are resampled to $25\mathrm{Hz}$ to match the robot orchestration rate. Vertex sequences are computed by subtracting the zero-pose subject-specific face meshes, then added to the designed human face template, $\mathbf{T}^{H}$, to get subject-agnostic sequences. Sequences are split $80/10/10$ into train/validation/test sets, with two subjects reserved for testing generalizability.
The model is implemented in PyTorch. The sliding window size $K$ is $8$ ($320\mathrm{ms}$). The hidden dimension $H$ is $64$. The robot has $U=4792$ vertices and $B=51$ blendshapes. Training uses the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$, learning rate $1e^{-4}$, weight decay $1e^{-4}$. $w_{m}$ is set to $1$. A dropout layer with a dropout rate of $0.1$ is added after each FC layer except the one with Sigmoid activation. $200$ epochs are trained on an NVIDIA 4090 GPU ($\sim 3\mathrm{sec}/\mathrm{epoch}$). Wav2vec2-XLSR53 is fixed during training. For inference, an $7\mathrm{Hz}$ $5$th-order low-pass Butterworth filter is applied. The model generates blendshape coefficients from speech at $>4000\mathrm{Hz}$ on GPU.
@@ VI-A Robot Development Experiments
Figure 6: Tracking Performance Validation. MSE error distributions between simulated and physically actuated facial articulation sequences are presented using violin plots, box and whisker plots, and scattered points, with each point representing one frame. Evaluation landmarks are grouped by facial region. Ten realistic facial articulation sequences from different speakers with distinct speaking styles were evaluated. Mean median errors across the ten sequences for each facial region are $2.56\mathrm{mm}$ (eye), $3.39\mathrm{mm}$ (brow), $1.74\mathrm{mm}$ (nose), $3.08\mathrm{mm}$ (cheek), $3.86\mathrm{mm}$ (mouth), and $5.03\mathrm{mm}$ (jaw). The results demonstrate that the animatronic robot face achieves accurate tracking performance across various facial regions and speaking styles.
@@
Motion Space Validation: The first experiment quantitatively validates the physical realization of the designed LBS motion space defined by the robot blendshapes. Achieving the comprehensive set of blendshapes is essential for enabling expressive facial articulation. A VICON motion capture system is employed to track the physical skinning deformations. Reflective markers are attached to key control points and auxiliary skinning locations, corresponding to landmarks in the simulator. By individually actuating each blendshape and comparing the VICON marker positions with the simulated landmarks, the discrepancies can be measured using the mean squared error (MSE).
The results, presented in Fig. 5, demonstrate highly accurate realization of the designed motion space. Across various facial regions, the MSE errors are consistently remain at the millimeter scale. For instance, the mouth blendshapes exhibit a median error of $3.76\mathrm{mm}$, while the eye blendshapes show a median error of $2.41\mathrm{mm}$. These low errors confirm the successful translation of the designed blendshapes into physical skinning deformations, validating the effectiveness of the skinning-oriented design approach.
Tracking Performance Validation: The second experiment evaluates the animatronic platform’s dynamic tracking performance in real-time by assessing its ability to follow reference facial motions. Achieving responsive and accurate tracking is crucial for enabling lifelike and synchronous facial expressions. The experimental setup employs the same VICON system, with markers attached to the robot’s control points and auxiliary skinning locations. A diverse set of $10$ realistic facial articulation sequences from different speakers with distinct speaking styles from the training set, enhanced with random blinks, serve as reference signals. These sequences encompass a wide range of full facial dynamics, representative of the system’s intended operating conditions.
As demonstrated in Fig. 6, the animatronic platform exhibits remarkable tracking performance. The MSE errors between the reference signals and the tracked positions remain consistently low over time across all facial regions and speaking styles. Comparing these dynamically achieved errors with those achieved statically during the motion space validation experiment presented in Fig. 5, the errors for different facial regions are very similar, on the millimeter scale. These results validate the effectiveness of the proposed electro-mechanical design in enabling precise and responsive facial motion control.
@@ VI-B Imitation Learning Experiments
Both the ground-truth and generated robot skinning motions are rendered as videos without facial textures to ensure that the evaluation focuses solely on motion quality. Participants are asked to choose which video in each pair shows more natural motion matching the speech or exhibits similar quality. The display order of the pair is randomized to prevent selection bias. Participants must first pass a qualification test to ensure meaningful results.
In total, $12$ qualified participants evaluated $80$ video pairs three times each. The results show that participants preferred the generated motions in $29.2\%$ of the tests and the ground truth in $45.0\%$, while judging $25.8\%$ of the examples to exhibit similar quality from both methods. These findings validate the effectiveness of the motion synthesis model in generating robot skinning motions from speech that are sufficiently natural and expressive to satisfy human perception.
The overall performance, including the motion synthesis on the developed animatronic robot face, is demonstrated in the supplementary video. The developed system is capable of automatically generating appropriate and dynamic facial expressions from speech in real-time, indicating the validity of the proposed skinning-centric approach that tightly integrates embodiment design and motion synthesis.
@@ VII Conclusions and Future Works
This paper introduces a novel, skinning-centric approach to drive animatronic robot facial expressions from speech. By employing LBS as the core representation, tightly integrated innovations in embodiment design and motion synthesis are achieved. LBS serves as a guiding principle for the actuation topology, enables seamless human expression retargeting, and facilitates speech-driven facial motion generation. Experimental results conclusively demonstrate that the developed animatronic robot face successfully generates highly realistic facial expressions from speech automatically and in real-time, validating the effectiveness of the proposed approach.
Future research can build upon this work in two primary directions. First, a general robot facial muscular system design approach could be explored that drives any human-like robot face simply by replacing the customized skull and skinning components, facilitating easy fabrication and control of customized animatronic faces. Second, more advanced imitation learning methods could empower speech-synchronized robot facial expressions with controllable emotions to enable complex facial signals that support richer, more engaging human-robot interactions. As animatronic robot technology continues to mature, realistic robotic faces will likely become commonplace across entertainment, education, healthcare, and other facets of life. This pioneering skinning-based approach establishes a strong foundation for that future.

### v3

@@ V-A Model Architecture
The model architecture $f(\cdot)$ takes raw speech waveforms $\mathbf{s}$ and generates blendshape coefficients $\boldsymbol{\theta}$ representing skinning motions. It contains: (i) a frame-level speech encoder extracting embeddings via a transformer-based phoneme logit extractor and temporal fusion module, (ii) a speaking style encoder embedding conditioning vectors, and (iii) an LBS encoder projecting conditioned speech embeddings to blendshape coefficients in the range $[0,1]$. In this context, speaking style refers to the characteristic facial motion patterns exhibited by different speakers during speech production. These patterns encompass individual variations in articulation, facial muscle engagement, and expressive tendencies, which contribute to the unique visual signature of each speaker’s facial movements during speech.
Specifically, the phoneme extractor follows the state-of-the-art transformer-based self-supervised pre-trained speech model, Wav2vec2 [29], finetuned on the CommonVoice dataset [30] with $53$ languages and $392$ phoneme classes to enhance cross-language generalization. It outputs a phoneme logits sequence at $49\mathrm{Hz}$ followed by resampling to $25\mathrm{Hz}$ to match the robot orchestration rate. The temporal fusion module progressively fuses $K$-step neighborhood logits over a sliding window into frame-level speech embeddings, comprising $M=\log_{2}K$ stacked Resnet blocks with a kernel size of $3$, stride of $2$, and $H$ filters. The $N$-way speaking style vector accounts for cross-subject variations and is embedded with a fully-connected (FC) layer having $H$ hidden units. This component is crucial for capturing and reproducing individual speaking styles, enhancing the model’s ability to generate diverse and personalized facial expressions. The LBS encoder takes the late-fused conditioned speech embedding as input and outputs blendshape coefficients with two successive layers: an FC layer having $2H$ hidden units with a ReLU activation, and an FC layer with $H$ hidden units followed by a Sigmoid activation.
@@ V-B Training
Figure 5: Motion Space Validation. Actuated blendshape error for different facial regions (left figure): Color-coded skinning landmarks represent different facial regions for evaluation. The 3D face model shows color-coded landmarks for different facial areas. Error distributions between simulated and physically actuated blendshapes are visualized using violin plots, box plots, and scattered points. Each point represents a single blendshape, evaluated using region-specific landmarks. Median errors (in mm) are provided for each facial region, ranging from $1.76\mathrm{mm}$ (nose) to $8.63\mathrm{mm}$ (jaw). Qualitative comparison (right figure): Visual comparison of eight simulated (gray mesh) versus actuated (realistic skin) blendshapes are shown. Blendshapes (1)-(6) demonstrate high accuracy, while (7) mouth close and (8) jaw open highlight limitations in the current design, exhibiting maximum errors for their respective regions.
@@ V-C Inference
The robot LBS decoder, a crucial component in the inference branch as illustrated in Fig. 4, converts generated blendshape coefficients into skinning motions. Utilizing pre-designed robot blendshapes, this decoder operates with fixed weights established during the facial motion space design phase. While structurally analogous to the human LBS decoder, its weights are determined by applying designed robot’s blendshapes to vertex displacements. The non-trainable nature of this decoder ensures consistent interpretation of blendshape coefficients throughout inference.
During the inference process, the trained model first generates a sequence of blendshape coefficients from input speech. This sequence undergoes temporal smoothing via a low-pass Butterworth filter to enhance stability. The robot LBS decoder then processes this filtered sequence, computing the skinning motion that serves as a reference signal for the downstream kinematics simulator. This approach maintains a consistent mapping between blendshape coefficients and resulting skinning motions, preserving the integrity of the designed facial expression space throughout inference.
@@ V-D Implementation Details
The VOCASET dataset [24], comprising audio-3D scan pairs of English utterances, is utilized for training and testing. This dataset was selected due to its high-quality 3D facial motion captures, diverse subjects, and phonetically balanced utterances, making it suitable for learning speech-driven facial expressions. VOCASET contains $255$ unique sentences shared among $12$ subjects, with $480$ facial motion sequences captured at $60\mathrm{Hz}$, each lasting $3$ to $4$ seconds. The 3D face mesh consists of $5023$ vertices.
For this study, VOCASET motions are resampled to $25\mathrm{Hz}$ to align with the robot orchestration rate. Subject-agnostic sequences are derived by subtracting zero-pose subject-specific face meshes from vertex sequences and adding them to the designed neutral (zero-pose) human face, $\mathbf{T}^{H}$. The data is split $80/10/10$ into train/validation/test sets, with two subjects reserved for testing generalizability. While VOCASET provides a foundation for initial model validation, the limited subject count may constrain generalization. This study serves as a proof-of-concept for the proposed approach, with exploration of larger, diverse datasets to enhance model generalizability being a potential area for future research.
The model is implemented in PyTorch with the following parameters: sliding window size $K=8$ ($320\mathrm{ms}$), hidden dimension $H=64$, and robot specifications of $U=4792$ vertices and $B=51$ blendshapes. Training utilizes the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$, learning rate $1e^{-4}$, and weight decay $1e^{-4}$. The motion weight $w_{m}$ is set to $1$. Dropout layers with a rate of $0.1$ are inserted after each fully connected (FC) layer, except for the one with Sigmoid activation. The model is trained for $200$ epochs on an NVIDIA 4090 GPU, taking approximately $3\mathrm{sec}/\mathrm{epoch}$. The Wav2vec2-XLSR53 component remains fixed during training. For inference, a $7\mathrm{Hz}$ $5$th-order low-pass Butterworth filter is applied to smooth the output. The model demonstrates high-speed performance, generating blendshape coefficients from speech at a rate exceeding $4000\mathrm{Hz}$ on GPU.
@@ VI-A Robot Development Experiments
Figure 6: Tracking Performance Validation. MSE error distributions between simulated and physically actuated facial articulation sequences are presented using violin plots, box and whisker plots, and scattered points, with each point representing one frame. Evaluation landmarks are grouped by facial region. Ten realistic facial articulation sequences from different speakers with distinct speaking styles were evaluated. Mean median errors across the ten sequences for each facial region are $2.56\mathrm{mm}$ (eye), $3.39\mathrm{mm}$ (brow), $1.74\mathrm{mm}$ (nose), $3.08\mathrm{mm}$ (cheek), $3.86\mathrm{mm}$ (mouth), and $5.03\mathrm{mm}$ (jaw). This comprehensive visualization demonstrates the animatronic face’s ability to maintain consistent tracking performance across diverse speaking styles while highlighting region-specific variations in accuracy.
@@
Motion Space Validation: This experiment quantitatively validates the physical realization of the designed LBS motion space defined by the robot blendshapes. Achieving the comprehensive set of blendshapes is essential for enabling expressive facial articulation. A VICON motion capture system is employed to track the physical skinning deformations. Reflective markers are attached to key control points and auxiliary skinning locations, corresponding to landmarks in the simulator. By individually actuating each blendshape and comparing the VICON marker positions with the simulated landmarks, the discrepancies can be measured using the mean squared error (MSE).
The results, presented in Fig. 5, demonstrate highly accurate realization of the designed motion space. Across various facial regions, the MSE errors consistently remain at the millimeter scale. For instance, the mouth blendshapes exhibit a median error of $3.76\mathrm{mm}$, while the eye blendshapes show a median error of $2.41\mathrm{mm}$. These low errors confirm the successful translation of the designed blendshapes into physical skinning deformations, validating the effectiveness of the skinning-oriented design approach.
Tracking Performance Validation: This experiment evaluates the animatronic platform’s dynamic tracking performance in real-time by assessing its ability to follow reference facial motions. Achieving responsive and accurate tracking is crucial for enabling lifelike and synchronous facial expressions. The experimental setup employs the same VICON system, with markers attached to the robot’s control points and auxiliary skinning locations. A diverse set of $10$ realistic facial articulation sequences from different speakers with distinct speaking styles from the training set, enhanced with random blinks, serve as reference signals. These sequences were carefully selected to encompass a wide range of full facial dynamics, representative of the system’s intended operating conditions and potential real-world applications.
As demonstrated in Fig. 6, the animatronic platform exhibits remarkable tracking performance. The MSE errors between the reference signals and the tracked positions remain consistently low over time across all facial regions and speaking styles. Comparing these dynamically achieved errors with those obtained statically during the Motion Space Validation experiment presented in Fig. 5, the errors for different facial regions are very similar, on the millimeter scale. These results validate the effectiveness of the proposed electro-mechanical design in enabling precise and responsive facial motion control, even under dynamic conditions mimicking natural speech production.
@@ VI-B Imitation Learning Experiments
Both ground-truth and generated motions are rendered as textureless videos to focus the evaluation solely on motion quality. Participants are asked to choose the more natural motion matching the speech or indicate similar quality. The display order is randomized to prevent bias. Participants must pass a qualification test to ensure meaningful results.
In total, $12$ qualified participants evaluated $80$ video pairs three times each. Results show participants preferred generated motions in $29.2\%$ of tests, ground truth in $45.0\%$, and similar quality in $25.8\%$. These findings validate the motion synthesis model’s effectiveness in generating natural, expressive robot skinning motions from speech, satisfying human perception.
The overall performance, including the motion synthesis on the developed animatronic robot face, is demonstrated in the supplementary video. The developed system is capable of automatically generating appropriate and dynamic facial expressions from speech in real-time, indicating the validity of the proposed skinning-centric approach that tightly integrates embodiment design and motion synthesis. However, it’s crucial to acknowledge the study’s limitations, such as the lack of in-person interactions with the physical robot and the limited participant pool. Future work should address these constraints to further validate the system’s effectiveness in real-world scenarios.
@@ VII Conclusions and Future Works
This paper introduces a novel, skinning-centric approach to drive animatronic robot facial expressions from speech. By employing LBS as the core representation, tightly integrated innovations in embodiment design and motion synthesis are achieved. LBS serves as a guiding principle for the actuation topology, enables seamless human expression retargeting, and facilitates speech-driven facial motion generation. Key technical contributions include the skinning-oriented design approach, the electro-mechanical implementation achieving millimeter-scale accuracy, and the imitation learning method for motion synthesis. Experimental results conclusively demonstrate that the developed animatronic robot face successfully generates highly realistic facial expressions from speech automatically and in real-time, validating the effectiveness of the proposed approach.
Future research can build upon this work in two primary directions. First, a general robot facial muscular system could be explored that drives any human-like robot face simply by replacing the customized skull and skinning components, facilitating easy fabrication and control of customized animatronic faces. This could involve developing modular actuation units and adaptive control algorithms to accommodate various facial structures. Second, more advanced imitation learning methods could empower speech-synchronized robot facial expressions with controllable emotions to enable complex facial signals that support richer, more engaging human-robot interactions. This might entail incorporating multi-modal inputs beyond speech and developing more sophisticated emotion models. As animatronic robot technology continues to mature, realistic robotic faces will likely become commonplace across entertainment, education, healthcare, and other facets of life. This pioneering skinning-based approach establishes a strong foundation for that future.
Acknowledgement: This work was supported in part by the National Natural Science Foundation of China (No.62376031).